HIGH RESOLUTION ELECTRONIC SPECTROSCOPY OF 2,6-DIAMINOPYRIDINE IN THE GAS PHASE^{a}

\underline{CASEY L. CLEMENTS}, ADAM J. FLEISHER, JUSTIN W. YOUNG, JESSICA A. THOMAS, AND DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260.

Ab initio calculations suggest that 2,6-diaminopyridine (26DAP) has several interesting low-frequency vibrations arising from motion of its amino groups. For this reason, 26DAP has been studied in the gas phase using both low resolution and high resolution electronic spectroscopy techniques. Presented here are the results of this study, which provide information about the structural and dynamical properties of 26DAP in both the ground and excited electronic states. The results will be discussed.

aWork supported by NSF (CHE-0615755).